Khaling Company sold 26,900 units last year at $16.50 each. Variable cost was $11.60, and total fixed cost was $136,710. Required:

1. Prepare an income statement for Khaling for last year.

2. Calculate the break-even point in units.

3. Calculate the units that Khaling must sell to earn operating income of $15,680 this year.

Required:
1. Prepare an income statement for Khaling for last year.
2. Calculate the break-even point in units.
3. Calculate the units that Khaling must sell to earn operating income of $15,680 this year.

Contribution Margin Income Statement

1. Prepare an income statement for Khaling for last year. Refer to the list of Amount Descriptions for the exact wording of text items within your income statement. If a net loss is incurred, enter that amount as a negative number using a minus sign.

Khaling Company

Contribution Margin Income Statement

For the Last Year

1

2

3

4

5

X

Unit Calculations

2. Calculate the break-even point in units.

3. Calculate the units that Khaling must sell to earn operating income of $15,680 this year.
Answer #1

1.

Sales Revenue $           4,43,850 =26900*16.5
Variable Expenses $           3,12,040 =26900*11.6
Contribution Margin $           1,31,810
Fixed Expenses $           1,36,710
Net Operating Income $               -4,900

2. Contribution margin per unit = $16.50 - 11.60 = $4.90

Break even units = Fixed costs / Contribution margin per unit
= $136710 / 4.90 = 27900 units

3.

Units to earn $15680 = (Fixed costs+Target Income) / Contribution margin per unit
= ($136710+15680) / 4.90 = 31100 units